The simplest rationalizing factor of 4th root of 48 is

Grade:8

1 Answers

Arun
5 years ago
48 = 2 × 2 × 2 × 2 × 3

4 th root of 48 = ( 2⁴ × 3 )^ 1/4 = 2 × 3^1/4

simplest rational factor = 3^ 3/4

2 × 3^ 1/4 × 3^3/4 = 2 × 3 = 6 ( rational number)

I hope this helps you.